OpenStudy (anonymous):

How much compound interest will an account of $10,000 bear after 5 years at 3% APR, compounded annually?

OpenStudy (anonymous.):

Do you know the formula to this ?

OpenStudy (anonymous.):

M=P(1+(r)/(n))^(n*t)

OpenStudy (anonymous.):

p=10,000 r=0.003 n=1 t=5

OpenStudy (anonymous.):

i=0.03 I mean :3

OpenStudy (anonymous):

M=10000(1+(0.03)/(1))^(1*5)

OpenStudy (anonymous.):

http://www.wolframalpha.com/ <--- Copy n Paste into here

OpenStudy (anonymous):

M=10000(1.03)^(5) M= 1.159 Right?

OpenStudy (anonymous):

Wait.....

OpenStudy (anonymous):

11592.7

OpenStudy (anonymous.):

Does that match any of your answer choices ?

OpenStudy (anonymous):

It does

OpenStudy (anonymous.):

:D !! there you go !! are taking a financial algebra class ?

OpenStudy (anonymous):

Thank you! It's called Buisness Math.. So yes? Haha